40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 2 01:47:28 2010, Alberto, Kiwamu, Update, IOO, mode measurement of new input optics afterMMT1.pngbeam_profile.pngMMT1_.dat.zipinside_vac_2.png
    Reply  Wed Jun 2 03:50:03 2010, Koji, Update, IOO, mode measurement of new input optics 
       Reply  Thu Jun 3 13:14:27 2010, kiwamu, Update, IOO, mode measurement of new input optics inside_vac_2.png
          Reply  Thu Jun 3 14:01:51 2010, kiwamu, Update, IOO, mode measurement of new input optics check_measurement_edit.png
             Reply  Thu Jun 3 14:13:20 2010, Jenne, Update, IOO, mode measurement of new input optics 
    Reply  Thu Jun 3 14:40:28 2010, Alberto, Kiwamu, Update, IOO, mode measurement of new input optics 
       Reply  Tue Jun 8 00:38:22 2010, Koji, Kiwamu, Update, IOO, improved Gaussian beam in new IOO 2010-6-7_2.png
          Reply  Fri Jun 11 14:31:04 2010, kiwamu, Update, IOO, mode matching of new IOO profile_MMT.pngdist_MMO.png
Message ID: 3054     Entry time: Tue Jun 8 00:38:22 2010     In reply to: 3046     Reply to this: 3068
Author: Koji, Kiwamu 
Type: Update 
Category: IOO 
Subject: improved Gaussian beam in new IOO 

The shape of the beam spot in the new input optics got much much better 

As Alberto and Kiwamu found on the last week, the beam spot after MMT1 had not been good. So far we postponed the mode measurement due to this bad beam profile.

Today after we did several things in the vacuum chamber, the beam spot became really a good Gaussian spot. See the attachment below.

There were two problems which had caused the bad profile:

(1)  a steering mirror after MMT1 with the incident angle of non 45 deg

(2) clipping at the Faraday.

 

Also MCT_QPD and MCT_CCD were recovered from misalignment  

Tomorrow we are going to restart the mode matching. 

 


(what we did)

* We started from checking the shape of the beam going out from the BS chamber. There still were some stripes which looked like an interference on the spot. 

* We found a steering mirror after MMT1 had the incident angle of non 45 deg. In fact the mirror had a large transmission. After we made the angle roughly 45 deg, the stripes disappeared.

However the spot still didn't look a good Gaussian, it looked slightly having a bump on the horizontal profile.

* Prior to moving of some optics in the vacuum, we ran the A2L_MC scripts in order to check the beam axis. And it was okay.

* To recover the MCT, we steered one of the vacuum mirrors which was located after the pick off mirror.  And after aligning some optics on the AP table, finally we got MCT recovered.

 * We rearranged MC_refl mirrors according to the new optical layout that Koji has made. At the same time the mirrors for IFO_refl was also rearranged coarsely.

 * We leveled the optical table of the MC chamber by moving some weights. Then we locked the MC again and aligned it. We again confirmed that the beam axis was still fine by running the A2L scripts.

 * We found the beam going through Faraday was off-centered by ~5mm toward the west. So we moved it so that the beam propagates on the center of it. 

 * Then looking at the beam profile after MMT1, we found that the profile became really nicer. It showed a beautiful Gaussian. 

In the attachment below, the top panel represents the horizontal profile and the bottom one represents the vertical profile.

The blue curves overlaid on the plot are fitted Gaussian profile, showing beautiful agreements with the measured profile.

Attachment 1: 2010-6-7_2.png  22 kB  | Hide | Hide all
2010-6-7_2.png
ELOG V3.1.3-